OpenStudy (anonymous):

A 25 ft ladder is leaning against a vertical wall. The bottom of the ladder is pulled horizontally away from the wall at 3 ft/sec. Determine how fast the top of the ladder is sliding when the bottom of the ladder is 15 ft from the wall <----- PLEASE EXPLAIN STEPS!

OpenStudy (anonymous):

 \[y=\sqrt{625-x ^{2}}\] \[dy/dt = (1/2)(-2x)/\sqrt{625-x ^{2}}(dx/dt)\] so at dx/dt =3 and x=15 dy/dt= -3/4 ft/s negative sign shows ladder is sliding downwards

OpenStudy (anonymous):

sorry multyply 3 also i forget to multiply 3...its -9/4 ft/s